What is y-10=-2(x-3) in standard form
Found 2 solutions by jim_thompson5910, lynnlo:
Answer by jim_thompson5910(35256)

y-10=-2(x-3)

y-10=-2x+6

y=-2x+6+10

y=-2x+16

2x+y=16

So the answer is 2x+y=16

Answer by lynnlo(4176)